# Vision-LSTM: xLSTM as Generic Vision Backbone

### v2

@@ 3.5 Classification Design
In order to perform classification with a ViT, the sequence of tokens is pooled into a single token, which is then used as input to a classification head. The most common methods to do this pooling are (i) adding a learnable [CLS] token at the start of the sequence or (ii) averaging all patch tokens to produce an [AVG] token. Whether to use the [CLS] or [AVG] token is typically a hyperparameter, where both variants achieve comparable performances. On the contrary, autoregressive models often require specialized classification designs. For example, Vim [67] requires the [CLS] token to be in the middle of the sequence, suffering heavy performance losses if other classification designs, e.g., an [AVG] token or two [CLS] tokens at start and end of the sequence, are employed. Due to its autoregressive nature, we explore different classification designs for ViL in Table 5. [AVG] averages all patch tokens, “Middle Patch” uses the middle patch token, “Middle [CLS]” uses a [CLS] token in the middle of the sequence, “Bilateral [AVG]” uses the average of the first and last patch token. We find ViL to be relatively robust the classification design where all performances that use a single token are within 0.6%. We only add a [CLS] token if it is used for classification, hence the best design (“Bilateral Concat”) does not use a [CLS] token.
Design Dim ACC
[AVG] 192 75.4
Middle Patch 192 75.9
Middle [CLS] 192 75.3
Bilateral [AVG] 192 75.9
Bilateral Concat 384 76.4
Table 5: How to pool the sequence of tokens for classification? ViL is relatively robust to the classification design. Using the middle patch token and averaging the first and last token (“Bilateral [AVG]”) work similarly well. The good performance of middle patch could be caused by the bias in ImageNet-1K that objects are typically in the center of the image. We therefore choose a bilateral design to avoid exploiting this bias. Additionally, concatenating the first and last token (“Bilateral Concat”) instead of averaging improves performance and is used as final design for ViL. We conduct this study by training a ViL-T on ImageNet-1K for 400 epochs.
3.6 Vision Adjustments of xLSTM
The xLSTM [5] architecture was designed for language modeling and therefore includes some design choices that are specific to language modeling. One of these design choices is a causal 1D convolution (with kernel size 4) to encode local context into the query and key input of the mLSTM layer. As images are neither causal nor 1D structures, we replace the causal 1D convolution with a 2D convolution (with kernel size 3). When training a ViL-T for 400 epochs, this improves ImageNet-1K accuracy from 76.1% to 76.7%.
Another language specific design choice is the absence of bias parameters in projection and layernorm [4] layers which improves training stability for larger models [12, 16]. As vision models are typically much smaller than language models, ViTs commonly use all bias parameters. When adding bias parameters, a ViL-T trained for 400 epochs improves from 76.7% to 77.1% on ImageNet-1K.
@@ 4 Conclusion
Motivated by the success of xLSTM in language modeling, we introduced ViL, an adaption of the xLSTM architecture to vision tasks. ViL processes a sequence of patch tokens in alternating fashion. Odd blocks process image patches row-wise from top left to bottom right and even blocks go from bottom right to top left. Our new architecture outperforms SSM-based vision architectures and also optimized ViT models on ImageNet-1K classification, VTAB-1K transfer classification and ADE20K semantic segmentation. Remarkably, ViL is able to outperform ViT training pipelines, which are the result of years of hyperparameter tuning and transformer improvements.
In the future, we see potential in applying ViL when high-resolution images are needed for optimal performance, such as semantic segmentation or medical imaging. In these settings, transformers suffer from high computational costs due to the quadratic complexity of self-attention, where ViL can use a chunked form to trade-off between its parallel form (quadratic complexity) and recurrent form (linear complexity). Additionally, improving pre-training schemes (e.g. via self-supervised learning), exploring better hyperparameter settings or investigating hierarchical architectures are promising future directions that could improve ViL even further.
@@ Acknowledgments
Appendix A Changelog
A.1 Version 1
In the first version of this paper, we trained models with a causal 1D convolution in the mLSTM layer (instead of a 2D convolution), “Bilateral [AVG]” pooling (instead of “Bilateral Concat”) and a constant resolution of 224 (instead of pre-training at 192 followed by a short fine-tuning on resolution 224). This version of the architecture is also used to conduct the studies in Table 4 and Table 5. For completeness, we report the accuracies of these models in Table 6. In the first version, we trained slightly deeper models (ViL-T+ with 29 blocks and ViL-S+ with 26 blocks) to match the parameter count of Vim-T and Vim-S. However, as the models with a regular depth of 24 were already better than Vim with less parameters, we omit them in future versions.
Method #Params ACC
ViL-T 6M 77.3
ViL-T+ 7M 78.1
ViL-S 23M 80.7
ViL-S+ 26M 80.9
ViL-B 88M 81.6
Model #Params ACC
Long-sequence fine-tuning
ViL-T+ 7M 80.0
ViL-S+ 26M 82.2
ViL-B 88M 82.7
Table 6: ImageNet-1K accuracies of models trained in the first version of this paper.
A.2 Version 2
The current version of the paper improves performance by (i) replacing the causal 1D convolution with a 2D convolution (with kernel size 3 and padding 1) (ii) using “Bilateral Concat” instead of “Bilateral [AVG]” pooling (iii) using bias parameters in projection and LayerNorm [4] layers (iv) pre-training models on 192x192 followed by a short fine-tuning on 224x224 resolution.
Appendix B Extended Results
B.1 Runtime Comparison of ViL vs Vim
We compare the time it takes to train ViL and Vim [67] for a single ImageNet-1K epoch in Table 7. We follow the scaling procedure of ViTs, using 192 (T), 384 (S), 768 (B), 1024 (L) as latent dimension where the large scale additionally doubles the number of blocks.
@@ B.1 Runtime Comparison of ViL vs Vim
Table 7: Runtime comparisons between Vim [67] and ViL. ViL is up to 69% faster despite the current lack of a custom CUDA kernel. As mLSTM (and ViL) can be parallelized analogous to FlashAttention [15, 14] via custom hardware optimizations, ViL will become even faster in the future. Runtimes denote the training time for 10 ImageNet-1K epochs and are extrapolated from short benchmark runs on a single A100-80GB-PCIe using float16 precision and 224x224 images.
B.2 Impact of Longer Training
We investigate the impact of training for a longer duration in Table 8. We also attempted to train larger models for longer but did not observe similar performance gains. One could probably change that by tuning hyperparameters for an 800 epoch training schedule, which we omit due to the large computational costs associated with training larger models.
@@ B.2 Impact of Longer Training
Table 8: Performance comparison of tiny models when training for 400 and 800 epochs. ADE20K mIoU uses single-scale evaluation. All settings follow the ones used in the main paper.
B.3 Long-sequence fine-tuning
Table 9 shows results for the long-sequence fine-tuning setting of Vim [67].
Method #Params ACC
DeiT-III-T (reimpl.) 6M 78.3
Vim-T [67] 7M 78.3
ViL-T 6M 80.1
DeiT-S [51] 22M 81.4
DeiT-III-S [54] 22M 83.2
Vim-S [67] 26M 81.6
ViL-S 23M 82.7
DeiT-B [51] 86M 83.2
DeiT-III-B [54] 86M 84.5
ViL-B 89M 83.3
Table 9: Fine-tuning ImageNet-1K pre-trained models by using a 50% overlap when creating patches. This setting increases the number of patch tokens from 196 to 729. ViL easily outperforms Vim but falls short of ViTs on larger models. Note that ViL FLOPS are calculated using the parallel form of the mLSTM, which scales quadratically with the sequence length. Using the chunkwise form instead of the parallel form can reduce FLOPS for longer sequences (see Section C.2).
B.4 VTAB-1K Individual Dataset Results
Table 10 presents accuracies for each individual dataset of the VTAB-1K [64] benchmark.
@@ B.5 Robustness
Appendix C Implementation Details
C.1 Hardware
We train models on a mix of custom hardware servers (mainly A100 and A40 GPUs) and public research clusters equiped with either 8xA100 or 4xA100 nodes.
@@ C.1 Hardware
C.2 FLOPS Calculation
We calculate the FLOPS with the fvcore library [^1] https://github.com/facebookresearch/fvcore. For ViL, we report the FLOPS of the mLSTM parallel form. Given sequence length $L$ and hidden dimension $D$, it has complexity $\mathcal{O}(L\frac{L}{2}D)$ where the $\frac{L}{2}$ stems from the causal masking.
We report FLOPS for a complexity of $\mathcal{O}(L\frac{L}{2}D)$ even though our current implementation of the mLSTM has complexity $\mathcal{O}(L^{2}D)$ (i.e., entries of the $QK$ matrix that are later masked are still calculated) which we justify by the fact that FlashAttention-2 [14] is approximately 1.7x faster with a causal mask than without. Therefore, an optimized hardware implementation of the mLSTM could also omit the calculation of values that are not needed.
@@ C.2 FLOPS Calculation
For the total pre-training compute in Figure 2, we consider an efficient implementation of stochastic depth [32] which omits the calculation of a dropped block instead of masking it. Therefore, we change the implementation of ViT [18] to use our efficient stochastic depth implementation. Vim does not use stochastic depth for training as their models are relatively small.
C.3 ViL Hyperparameters
@@
   Linear LR Scaling Divisor [22] 1024
@@
Precision mixed bfloat16 [35]
@@
   Mixup [65] $\alpha$ 0.8
   Cutmix [63] $\alpha$ 1.0
@@
C.4 Fine-tuning on VTAB-1K
@@
   Layer-wise lr decay [13] 0.65*
@@
Precision mixed bfloat16 [35]
@@
C.5 ADE20K Semantic Segmentation Fine-tuning
@@
   Linear LR Scaling Divisor [22] 16
   Layer-wise lr decay [13] 0.65*
@@
C.6 DeiT-III Reimplementation Hyperparameters
Table 12 shows detailed hyperparameters used to train DeiT-III-T (reimpl.) from Table 1.
@@ C.6 DeiT-III Reimplementation Hyperparameters
Stochastic depth [32]
Layerscale [52] 1e-4
Optimizer AdamW [42, 36]
@@
   Linear LR Scaling Divisor [22] 1024
@@
Precision mixed bfloat16 [35]
@@
   Mixup [65] $\alpha$ 0.8
   Cutmix [63] $\alpha$ 1.0

### v3

@@ 4.2 Classification Design
In order to perform classification from a sequence of tokens, it is common to aggregate information from the whole sequence, which is then used as input to a classification head. The most common methods to do this aggregation are (i) adding a learnable [CLS] token to the input sequence or (ii) averaging all patch tokens to produce an [AVG] token. In ViTs, whether to use the [CLS] or [AVG] token is typically a hyperparameter, where both variants achieve comparable performances. On the contrary, other sequence models models often require specialized classification designs. For example, Vim [82] requires the [CLS] token to be in the middle of the sequence, suffering heavy performance losses if other classification designs, e.g., an [AVG] token or two [CLS] tokens at start and end of the sequence, are employed.
We explore different classification designs for ViL in Table 5. (a) We choose concatenating the first and last patch as aggregation method due to its strong classification performance. As our final models also perform well in semantic segmentation (see Table 2), we do not retrain models with [AVG] aggregation even though the ablation suggests that this could boost performance even further for segmentation tasks. (b) Adding learnable [CLS] tokens show no benefit. Therefore, we do not use any [CLS] tokens for ViL.
Aggregation IN1K ADE20K
Bilateral Mean 73.0 31.5
Bilateral Concat 73.7 31.7
[[[[AVG]]]] 72.6 32.8
Center [AVG] 72.4 32.1
Aggregation IN1K
Concat Bilateral Patches 73.7
Mid [CLS] 71.8
Bilateral [CLS] 73.5
Mid + Bilateral [CLS] 73.0
Table 5: Classification design. (a) ViL aggregates classification information well in the first and the last patches (bilateral), leading to good classification performance if the first and last patches are averaged or concatenated. Averaging all patches ([AVG]) or the 4 center patches (Center [AVG]) results in strong segmentation performances but lackluster classification performances. (b) Adding learnable [CLS] tokens to the start and end of the input sequence (Bilateral [CLS]) offers no benefit over simply using the first and the last patch. Incorporating a [CLS] token in the middle of the sequence, akin to Vim [82], does not improve performance. Default settings
5 Limitations and Future Work
The biggest limitation of ViL is the current lack of an optimized hardware implementation of the mLSTM, which results in longer runtimes than ViTs, which have multiple optimized hardware implementations [18, 17]. This makes a runtime/throughput analysis of models, a vital metric to judge practicability, difficult as the practical relevance of inefficient implementations is quite low. As a proxy, we report FLOP counts, where ViL is comparable to ViT on low-resolution tasks and far better than ViT on high-resolution tasks due to its linear complexity. While FLOPS are far from an optimal proxy for runtime/throughput, they suggest that ViL can be much faster than ViT on high-resolution tasks once an optimized hardware implementation exists. Note that ViL is already faster than Vim (see Appendix A.1) despite its optimized hardware implementation.
This limitation snowballs in multiple other directions. For example, scaling model size further, tuning hyperparameters, training on larger datasets, exploring self-supervised pre-training or investigating hierarchical architectures are all interesting avenues for future work that are currently quite costly due to the lack of an optimized hardware implementation.
Please note that this is merely a technical limitation, not a methodical one as the mLSTM is heavily parallelizable. However, implementing fast compute kernels in CUDA [54] or Triton [63] is highly non-trivial as it requires expert hardware architecture knowledge, advanced implementation skills and potentially multiple development cycles to iron out numerical inaccuracies or instabilities.
However, the results of recent linear attention mechanisms show impressive FLOPS utilization (e.g., [78]). As the mLSTM can be parallelized with similar techniques it is only a matter of time that the mLSTM achieves a similar FLOPS utilization, which will make the mLSTM faster than transformers once an efficient hardware implementation is available.
Additionally, we made a significant effort to make our architecture as efficient as possible, using the tools that are currently available to us. Notably, our architecture is already much faster (up to 70%) than Vim [82] despite Vim using a custom CUDA kernel, as shown in Appendix A.1. For reference, in language modeling, Mamba is roughly on-par with transformers in terms of speed and 4x faster than than the xLSTM (as mentioned in [5]), again, due to the current lack of efficient hardware implementation of the mLSTM. These considerations further underline the potential of our simple and efficient design for vision applications.
6 Related Work
Generic Vision Backbones.
The inductive bias of CNNs [25, 47] has demonstrated ground-breaking advancements in computer vision [46] in the early deep learning days. Features of CNNs have been found to learn generic visual features that can be used for a variety of tasks [21]. Subsequently, countless works improved various aspects such as architectures [60, 33, 41, 61, 50] or pre-training strategy [20, 53, 80, 27, 12, 28].
Sequence Models in Vision.
The introduction of transformers [70] demonstrated exceptional scalability in language processing, which motivated the vision community to explore transformers also in computer vision [11, 16] but was applied on pixels or small patches which inhibited large costs due to the quadratic complexity of self-attention. This restriction was alleviated by the seminal work Vision Transformers (ViTs) [22] by using larger patches to aggregate local information and reduce training costs. Similar to CNNs, lots of work improved on the ViT architecture by refining training procedures [64, 65, 67, 9, 4, 76, 34]. The recent advancement of autoregressive models in language processing [30, 58] has also gathered interest in the vision community [82, 23] due to the linear scaling property which allows applications to high-resolution tasks such as medical imaging [51] or video understanding [48].
@@ 7 Conclusion
Motivated by the success of xLSTM in language modeling, we introduced ViL, an adaption of the xLSTM architecture to vision tasks. ViL processes a sequence of patch tokens in alternating fashion. Odd blocks process image patches row-wise from top left to bottom right and even blocks go row-wise from bottom right to top left. Our new architecture outperforms SSM-based vision architectures, other autoregressive vision architectures and also optimized ViT models on ImageNet-1K classification, VTAB-1K transfer classification and ADE20K semantic segmentation. Remarkably, ViL is able to outperform ViT training pipelines, which are the result of years of hyperparameter tuning and transformer improvements.
In the future, we see potential in applying ViL when high-resolution images are needed for optimal performance, such as semantic segmentation or medical imaging. In these settings, transformers suffer from high computational costs due to the quadratic complexity of self-attention, where the linear complexity of ViL allows compute efficient processing of long sequences. Additionally, improving pre-training schemes (e.g., via self-supervised learning), exploring better hyperparameter settings or investigating hierarchical architectures are promising future directions.
@@ Acknowledgments
Appendix A Extended Results
A.1 Runtime Comparison of ViL vs Vim
We compare the runtime to train ViL and Vim [82] for 10 ImageNet-1K epochs in Table 6. We follow the scaling procedure of ViTs, using 192 (T), 384 (S), 768 (B), 1024 (L) as hidden dimension where the (L)arge scale doubles the number of blocks.
@@ A.1 Runtime Comparison of ViL vs Vim
Table 6: Runtime comparisons between Vim [82] and ViL. ViL is up to 69% faster despite the current lack of a optimized hardware implementation. As mLSTM (and ViL) can be parallelized analogous to FlashAttention [18, 17] via custom hardware optimizations, ViL will become even faster in the future. Runtimes denote the training time for 10 ImageNet-1K epochs and are extrapolated from short benchmark runs on a single A100-80GB-PCIe using float16 precision and 224x224 images.
A.2 Impact of Longer Training
We investigate the impact of training for a longer duration in Table 7.
@@ A.2 Impact of Longer Training
Table 7: Performance comparison of tiny models trained for 400 and 800 epochs. ADE20K mIoU uses single-scale evaluation. All settings follow the ones used in the main paper.
A.3 VTAB-1K Individual Dataset Results
Table 8 presents accuracies for each individual dataset of the VTAB-1K benchmark.
@@ A.4 Robustness and Domain Generalization
Appendix B Implementation Details
B.1 Hardware
We train models on servers with either 8xA100 or 4xA100 nodes.
@@ B.1 Hardware
B.2 FLOPS Calculation
We use the fvcorehttps://github.com/facebookresearch/fvcore library to count FLOPS and report FLOPS of the mLSTM chunkwise form as described in Section 2.2. For the parallel parts, we report FLOPS for a complexity of $\mathcal{O}\big{(}(\frac{S}{2}+1)Sd\big{)}$ because the upper triangular entries of the $\mathbf{QK}$ matrix do not need to be calculated due to the causal structure. We justify this by the fact that FlashAttention-2 [17] is approximately 1.7x faster with a causal mask than without. Therefore, an optimized hardware implementation of the mLSTM could also omit the calculation of the upper triangular part of $\mathbf{QK}$.
@@ B.2 FLOPS Calculation
For the total pre-training compute in Figure 3, we consider an efficient implementation of stochastic depth [40, 68] which omits the calculation of a dropped block instead of masking it. Therefore, we change the implementation of ViT [22] to use our efficient stochastic depth implementation. Vim does not use stochastic depth for training as they only train tiny and small models.
B.3 ViL Hyperparameters
@@
   Linear LR Scaling Divisor 1024
@@
Precision mixed bfloat16
@@
   Mixup $\alpha$ 0.8
   Cutmix $\alpha$ 1.0
@@
B.4 Fine-tuning on VTAB-1K
@@
   Layer-wise lr deca 0.65*
@@
Precision mixed bfloat16
@@
B.5 ADE20K Semantic Segmentation Fine-tuning
@@
   Linear LR Scaling Divisor 16
   Layer-wise lr decay 0.65*
@@
B.6 DeiT-III Reimplementation Hyperparameters
Table 10 shows detailed hyperparameters used to train DeiT-III-T (reimpl.) from Table 1. Our reimplementation easily outperforms older baselines like DeiT-II-T (+2.7% ImageNet-1K accuracy) and is approximately even with the original on ADE20K (40.1 vs 39.8 on mIoU single-scale, 41.8 vs 42.2 mIoU multi-scale).
@@ B.6 DeiT-III Reimplementation Hyperparameters
Stochastic depth
Layerscale 1e-4
Optimizer AdamW
@@
   Linear LR Scaling Divisor 1024
@@
Precision mixed bfloat16
@@
   Mixup $\alpha$ 0.8
   Cutmix $\alpha$ 1.0